Microwave spectrum of the Xe–propyne Van der Waals complex

\underline{Th. Brupbacher}, W. Jäger^{a}, M. C. L. Gerry , Department of Chemistry, University of British Columbia, Vancouver, B.C., V6T 1Z1, Canada; D. F. Eggers †, Department of Chemistry, University of Washington, Seattle, WA 98195, U. S. A..

The microwave spectrum of the Xe-propyne Van der Waals complex has been recorded between 4—22 GHz, using a newly-developed pulsed molecular beam Fourier transform microwave spectrometer incorporating automated frequency scanning. The spectra of five isotopomers, involving ^{129}Xe, ^{131}Xe, ^{132}Xe, ^{134}Xe, and ^{136}Xe, were observed in natural abundance. The geometry of Xe-propyne is T-shaped, comparable to that of Ar – propyne^{b}. Hindered internal rotation of the methyl group in propyne splits the levels into A and E species. The A species transitions could be analysed using predictions from an on-going analysis of the infrared spectrum^{c}. The microwave spectrum was analysed using a rotation-internal rotation Hamiltonian. The barrier to internal rotation V_{3} was found to be ∼ 22cm^{−1}, comparable to that of Ar-propyne.

aPresent address: Department of Chemistry, University of Alberta, Edmonton, Alta., Canada, T6G 2G2

bT. A. Blake et al., J. Chem. Phys., 98, p. 6031 (1993)

cR. O. Watts, T A. Blake, P. Muino and D. F. Eggers †, private communication